OpenStudy (anonymous):

? HELP ?? Salesperson purchased car averaging 27 mi/gal ciry 38 mi/gal highway. Recent trip 1654 miles required 47 gallons. How many miles driven were city miles?

OpenStudy (anonymous):

if \(x\) miles were city, then \(1654-x\) miles were highway

OpenStudy (anonymous):

then you have \(\frac{x}{27}\) gallons were used in the city and \(\frac{1654-x}{38}\)gallons were used on the highway

OpenStudy (anonymous):

the total number of gallons used was \[\frac{x}{27}+\frac{1654-x}{38}=47\] solve for \(x\)

OpenStudy (anonymous):

if you don't want to do the algebra to solve for \(x\) you can cheat

OpenStudy (anonymous):

im confused

OpenStudy (anonymous):

me too

OpenStudy (anonymous):

there's alot of information and im trying to find the miles for the city

OpenStudy (anonymous):

did you understand the equation i wrote?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

ok the total number of miles driven is \(1654\) right?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

so if you drive 600 miles in the city, what is left for the highway miles?

OpenStudy (anonymous):

how did you get 600 miles

OpenStudy (anonymous):

i made it up it is not the answer, i am just asking if the total is 1654 and you drive 600 city, what is left for highway?

OpenStudy (anonymous):

1054

OpenStudy (anonymous):

do i find a common denominator or just multiply the denominators to 47

OpenStudy (anonymous):

ok so 1054 looks good suppose you drive \(x\) miles in the city, what is the highway miles?

OpenStudy (anonymous):

idk

OpenStudy (anonymous):

you answer has to have an \(x\) in it if you drive \(x\) miles in the city, then you drive \(1654-x\) miles on the highway

OpenStudy (anonymous):

ok i have x/27 + 1654/38 =47 and thats as ar as ive gotten do i mutiply the denominators to 47

OpenStudy (anonymous):

hello